# Decoding Secret Memorization in Code LLMs Through Token-Level Characterization

### v1

@@ III-A Motivation
While existing researches have shown that Code LLMs can memorize secrets or privacy from training datasets [37, 38, 39], the characteristics of these memorized secrets remain under-explored. Additionally, given that Code LLMs may generate fake secrets due to hallucination issues [8], it is necessary to explore the differences between the characteristics of real and fake secrets [40]. To this end, we conduct an in-depth analysis of secret memorization in Code LLMs at the token level. Following the prior study [8], we target at extracting 6 types of secrets that are listed in Table I. The secrets follow specific formats as described by the corresponding regular expressions in the table.
@@ III-B1 Source File Collection
For the 6 types of secrets, we use their regular expressions to search for code files containing them from open-source repositories through SourceGraph [41]. We choose SourceGraph for two reasons: (i) It integrates a vast number of open-source code repositories; (ii) It supports efficient code search using regular expressions, which perfectly meets our requirement for searching secrets.
In the search process, we only consider 5 languages, namely HTML, Java, JavaScript, PHP, and Python, which are commonly used in web applications that may require the 6 types of secrets for online API requests. After duplication, 31,978 code files are identified. To ensure data quality, we apply the filtering rules adopted in the StarCoder project [42] . Files that meet the following criteria are excluded:
@@
These filtering operations remove approximately 3.85% of the files. Subsequently, we perform an additional examination of the characters surrounding the matched content to filter out mismatched code files. Ultimately, 66.57% of the original search results are retained, resulting a set of 25,353 code files. Table II shows the language and secret distribution in the final results.
@@ III-B2 Completion Prompt Construction
For each code file, we locate the secret’s position and remove all content after the fixed secret prefix (e.g.AIza” for Google API Key). We then retain the line containing the secret and up to 7 preceding lines as the completion prompt. For example, as illustrated in Fig. 1, the fixed prefix “AIza” of a found Google API Key is retained, while all subsequent content is removed. After processing all code files, we obtain 2,400 completion prompts.
@@ III-B3 LLM-based Secret Generation
We feed these 2,400 prompts into a Proxy Code LLM to perform the code completion and predict the missing Google API Keys. In this study, we select StarCoder2-15B as the proxy LLM, which is a 15B-parameter model trained on multiple programming languages from The Stack v2 [43], supporting various tasks such as code completion and code generation.
StarCoder2-15B generates a token sequence step by step based on an input prompt. At each step, it generates a token probability distribution based on the context provided by the prompt and the previously predicted tokens. To avoid generating erroneous secrets that violate the format constraints in Table I, we set the probabilities of invalid tokens to 0. We then apply beam search with a width of 5 to select the next token, keeping the token sequence with the highest likelihood score as the final generated secret. During the generation process, we record the predicted token probability distribution of each step for further analysis.
@@ III-C Characterization
Based on our observation and analysis of token probabilities, we characterize the real and fake secrets generated by the Code LLM as follows. Note that these characteristics are generalizable to different types of secrets, we simply use Google API Keys to illustrate (see our replication package [45]).
C1: Tokens in real secrets tend to stabilize at relatively high probabilities after initial few decoding steps. Fig. 2 shows the scatter plot of token probabilities for both real and fake secrets (blue and red points respectively) along the decoding steps, where each point ($i$, $prob$) represents a token at the $i$-th decoding step with a probability of $prob$. Starting from the 5th step, the probability of most real secret tokens stabilizes at a relatively high level. For instance, at the 10th step, tokens of the 100 real secrets have probabilities over 0.9.
C2: Tokens in real secrets generally have higher probabilities than those in fake ones. According to Fig. 2, the overall probabilities of real secret tokens are generally higher than those of fake secret tokens, with the latter often more scattered between 0 and 1. For example, at the 10-th step, the probability values of many fake secret tokens are evenly distributed between 0 and 0.9.
C3: Probability advantage of tokens in real secrets is typically more pronounced than in fake ones. The probability advantage of a token at a given decoding step is defined as the difference between its probability and the probability of the next token in the distribution, ranked by descending probabilities. Fig. 3 shows a scatter plot of token probability advantages for both real and fake secrets along the decoding steps, where each point ($i$, $adv$) represents a token at the $i$-th decoding step with a probability advantage of $adv$. Similar to C2, the overall probability advantages of real secret tokens are generally higher than those of fake secret tokens.
C4: Certain fake secrets can be identified early in decoding process using secret strength metrics like Shannon entropy. Code LLMs often generate weak secrets containing continuously repeated tokens due to the decoding objective of maximizing likelihood [46]. As the repetition of characters or tokens increases during the decoding process of such fake secrets, the Shannon entropy of the generated token sequence continuously decreases. This observation inspires us to inspect the Shannon entropy during the decoding process and avoid certain fake secrets at an early stage to increase the likelihood of generating real secrets. For example, if the current decoding step token sequence is “SyA2-”, and “2” and “x” are the two tokens with the highest probabilities predicted by the Code LLM, we can discard “2” and choose “x” as the next token since appending “2” would decrease the sequence’s entropy, thus preventing Code LLMs from falling into token repetition and producing fake secrets like “AIzaSyA2-2-2-2...”.
@@ IV-A Token-Level Features
Average Probability ($avg\_prob$): According to C2, real secret tokens generally have higher probabilities than fake ones. To capture this while addressing the instability of using a single token’s probability, we use the average probability of tokens selected in previous decoding steps as a feature. In Fig. 5, the probabilities of tokens up to “2” are 0.99, 0.18, 0.03, 0.02, and 0.03, with an average of 0.25, which is set as the Average Probability for token “2”.
@@ IV-B Token Scoring Model Construction
Training Data Construction. We follow a process similar to our characterization study setup (Section III-B). We create completion prompts for each secret type based on searched code files and feed them into a Proxy Code LLM to generate candidate secrets. The generated candidates are validated and split into a real set $\mathcal{R}$ and a fake set $\mathcal{F}$. During the proxy Code LLM’s token decoding process, we extract the feature vector $\mathbf{feat}$ for each token, resulting in two vector sets: $\mathbf{Feat}\mathcal{R}$ and $\mathbf{Feat}\mathcal{F}$, corresponding to $\mathcal{R}$ and $\mathcal{F}$, respectively.
Scoring Model Training. To effectively distinguish between tokens of real and fake secrets, we train a Linear Discriminant Analysis (LDA) [47] model to find a combination of the four features that maximizes the differentiation between $\mathbf{Feat}\mathcal{R}$ and $\mathbf{Feat}\mathcal{F}$. We choose this linear model over complex models like Multi-Layer Perceptron (MLP) [48] because our feature vector for each token contains only four features, and using non-linear models can lead to overfitting issues with low-dimensional feature vectors [49].
@@ 10             $Beam\leftarrow Cands$
  • Masking Invalid Tokens (line 8). In traditional beam search, the LLM predicts the probability distribution $\mathbf{p}$ based on the existing token sequence $seq$ (line 7), and the next step is to select the top $B$ tokens from this distribution (line 9). However, in our secret extraction scenario, some tokens with unsupported characters are invalid (e.g., “*”), as secrets must adhere to specific formats. To address this, we mask invalid tokens in $\mathbf{p}$ by setting their probabilities to 0 (i.e. the MaskInvalidToks procedure), ensuring they are not selected in the subsequent step. The specific constraint for determining invalid tokens for each secret type is listed as a regular expression in Table I.
@@
    • Calling Scoring Model (line 12). Based on the extracted $\mathbf{feat}$, we call the trained scoring model to predict $prob_{\mathcal{R}}$ for $tok$ using Equation 1, indicating the probability that $tok$ belongs to a real secret.
@@
Additional Optimization. Based on C1 from our characterization study, we optimize the beam search process (lines 18-20) by retaining all expanded candidate hypotheses ($Cands$) during the first $K$ steps instead of selecting the top $B$ hypotheses. This prevents missing real secrets due to the initially low probabilities of early tokens [50]. We set $K$ to 4 based on our observations in C1 and computational cost considerations.
@@ V Experimental Setup
  • RQ3: What are the effects of components of DeSec on secret extraction?
  • RQ4: How effective is our token scoring model in identifying secret tokens?
V-A Implementation
In the online decoding, the beam search size $B$ is set to 5. The hyper-parameter w for the LLM-predicted token probability is set to 0.75 by evaluating the results within interval [0,1]. Fig. 6 shows that it yields the highest number of real secrets (1,076) and the best overall performance for generating both plausible secrets (PS#) and real secrets (RS#).
V-B Evaluation Dataset
@@
V-C Victim Code LLMs
@@ VI Results and Analysis
Comparison to HCR. DeSec significantly outperforms HCR in extracting plausible secrets, with PRimprovements ranging from 16.8% to 51.2% across the five victim Code LLMs. DeSec outperforms HCR for most secret types, except for GOCI in DeepSeek-Coder-6.7B and CodeLlama-13B, where HCR extracts 1 and 9 more plausible secrets, respectively. The superiority of DeSec can be attributed to: (i) Code LLMs generally performing better on completion tasks than filling tasks [54], corresponding to the prompt construction strategies in DeSec and HCR, and (ii) beam search decoding in DeSec being more effective than greedy search in HCR, evidenced by comparing HCR and BS-5 results.
@@
DeSec consistently demonstrates high effectiveness in extracting real secrets across all five victim Code LLMs, extracting between 95 and 358 real secrets. Interestingly, despite the scoring model being trained using StarCoder-15B as the proxy Code LLM, DeSec exhibits superior RS# for DeepSeek-Coder-6.7B (214) and CodeLlama-13B (358) compared to StarCoder-15B (210), suggesting that the identified characteristics (C1-C4) and the trained scoring model are generalizable to various Code LLMs. The higher RS# for CodeLlama-13B may indicate that it memorizes more real secrets, posing more severe privacy leakage risks.
@@
VI-C RQ3: Ablation Study
To investigate the effects of the two key enhancements in DeSec, namely masking invalid tokens and scoring model guided probability re-weighting, we conduct an ablation study. We respectively remove the two enhancements, resulting in two variants of DeSec: w/o masking and w/o scoring.
@@ VI-C RQ3: Ablation Study
Answer to RQ3 Token scoring model effectively prompts the models to output memorized secrets. Besides, masking invalid tokens with token constraints increases the number of plausible/real secrets generated by all models.
VI-D RQ4: Effectiveness in Identifying Secret Tokens
@@
Answer to RQ4 The token scoring model demonstrates high values across various metrics for most models, indicating its effectiveness in identifying secret tokens.
VI-E Case Study
Fig. 8 illustrates three secret generation examples for DeSec, HCR, and BS-5. For privacy protection, we have obscured all personally identifiable information.

### v2

@@ III-A Motivation
While existing researches have shown that Code LLMs can memorize secrets or privacy from training datasets [37, 38, 39], the characteristics of these memorized secrets remain under-explored. Additionally, given that Code LLMs may generate fake secrets due to hallucination issues [8], it is necessary to explore the differences between the characteristics of real and fake secrets [40].
To this end, we conduct an in-depth analysis of secret memorization in Code LLMs at the token level. We aim to extract the 6 most common types of commercial secrets included in the prior study [8], which together constitute 85.8% of all secrets. For ease of management, these secrets follow structured formats that can be defined using regular expressions. Table I lists these 6 types of secrets along with their corresponding regular expressions.
@@ III-B1 Source File Collection
For the 6 types of secrets, we use their regular expressions to search for code files containing them from open-source repositories through SourceGraph [41]. We choose SourceGraph for two reasons: (i) It integrates a vast number of open-source code repositories;
(ii) It supports efficient code search using regular expressions, which perfectly meets our requirement for searching secrets.
In the search process, we only consider 5 languages, namely HTML, Java, JavaScript, PHP, and Python, which are commonly used in web applications that may require the 6 types of secrets for online API requests. After duplication, 31,978 code files are identified as of June 16, 2024. To ensure data quality, we apply the filtering rules adopted in the StarCoder project [42] . Files that meet the following criteria are excluded:
@@
These filtering operations remove approximately 3.85% of the files. Subsequently, we perform an additional examination of the characters surrounding the matched content to filter out the mismatched code files. For example, the middle part (highlighted in blue) of “…kBdAIza…BVV…” matches the regular expression for Google API Key, but the surrounding characters indicate that it is not a valid Google API Key. Ultimately, 66.57% of the original search results are retained, resulting a set of 25,353 code files. Table II shows the language and secret distribution in the final results.
@@ III-B2 Completion Prompt Construction
For each code file, we locate the secret’s position and remove all content after the fixed secret prefix (e.g.AIza” for Google API Key). We then retain the line containing the secret and up to 7 preceding lines as the completion prompt, which provides sufficient context without overwhelming the model based on the preliminary observations and previous study [43]. For example, as illustrated in Fig. 1, the fixed prefix “AIza” of a found Google API Key is retained, while all subsequent content is removed. After processing all code files, we obtain 2,400 completion prompts.
@@ III-B3 LLM-based Secret Generation
We feed these 2,400 prompts into a Proxy Code LLM to perform the code completion and predict the missing Google API Keys. In this study, we select StarCoder2-15B as the proxy LLM, which is a 15B-parameter model trained on multiple programming languages from The Stack v2 [44], supporting various tasks such as code completion and code generation. StarCoder2-15B is chosen as the proxy LLM because it shares the same common decoder-only architecture as the other Code LLMs in our study. This ensures consistency in studying secret token generation.
StarCoder2-15B generates a token sequence step by step based on an input prompt. At each step, it generates a token probability distribution based on the context provided by the prompt and the previously predicted tokens. To avoid generating erroneous secrets that violate the format constraints in Table I, we set the probabilities of invalid tokens to 0. We then apply beam search with a width of 5 to select the next token, keeping the token sequence with the highest likelihood score as the final generated secret. The generation stops once the target secret’s required length (character number) is reached. During the generation process, we record the predicted token probability distribution of each step for further analysis.
@@ III-C Characterization
Based on our observation and analysis of token probabilities, we characterize the real and fake secrets generated by the Code LLM as follows. Note that these characteristics are generalizable to different types of secrets, we simply use Google API Keys to illustrate (see our replication package [46] for STSK, TCSI, and GOCI).
@@
C1: Tokens in real secrets tend to stabilize at relatively high probabilities after initial few decoding steps. Fig. 2 shows the scatter plot of token probabilities for both real and fake secrets (blue and red points respectively) along the decoding steps, where each point ($i$, $prob$) represents a token at the $i$-th decoding step with a probability of $prob$. Typically, the probabilities of the initial few steps (e.g., the first five) for real secret tokens are relatively low. However, as the decoding progresses, the probability values rapidly increase and stabilize at a relatively high level. For instance, at the 10th step, tokens of the 100 real secrets have probabilities over 0.9. This observation can be attributed to the model’s initial uncertainty in earlier steps; however, as it decodes more tokens and aligns with memorized sequences, its confidence increases quickly. Besides, an interesting discovery is that real secret tokens typically consist of multiple characters, while fake ones often have one single character. Consequently, real secrets decode in fewer steps and terminate earlier.
C2: Tokens in real secrets generally have higher probabilities than those in fake ones. According to Fig. 2, the overall probabilities of real secret tokens are generally higher than those of fake secret tokens, with the latter often more scattered between 0 and 1. For example, at the 10-th step, the probability values of many fake secret tokens are evenly distributed between 0 and 0.9. This can be explained by the internal memory mechanism of LLMs. In addition, we conducted t-tests for this characteristic, and the results indicate a significant probability difference between real tokens and fake tokens for all secret types ($p\ll 0.01$).
C3: Probability advantage of tokens in real secrets is typically more pronounced than in fake ones. The probability advantage of a token at a given decoding step is defined as the difference between its probability and the probability of the next token in the distribution, ranked by descending probabilities. Fig. 3 shows a scatter plot of token probability advantages for both real and fake secrets along the decoding steps, where each point ($i$, $adv$) represents a token at the $i$-th decoding step with a probability advantage of $adv$. Similar to C2, the overall probability advantages of real secret tokens are generally higher than those of fake secret tokens. In addition, we conducted t-tests for this characteristic, and the results indicate a significant probability advantage difference between real tokens and fake tokens for all secret types ($p\ll 0.01$).
C4: Certain fake secrets can be identified early in decoding process using secret strength metrics like Shannon entropy. Code LLMs often generate weak secrets containing continuously repeated tokens due to the decoding objective of maximizing likelihood [47]. As the repetition of characters or tokens increases during the decoding process of such fake secrets, the Shannon entropy of the generated token sequence continuously decreases. This observation inspires us to continuously inspect the Shannon entropy during the decoding process and avoid certain fake secrets at an early stage to increase the likelihood of generating real secrets. Although entropy is commonly used to assess the strength of secrets, we are the first to integrate it directly into the LLMs’ decoding process, rather than as a post-processing step. For example, if the current decoding step token sequence is “SyA2-”, and “2” and “x” are the two tokens with the highest probabilities predicted by the Code LLM, we can discard “2” and choose “x” as the next token since appending “2” would decrease the sequence’s entropy, thus preventing Code LLMs from falling into token repetition and producing fake secrets like “AIzaSyA2-2-2-2...”.
@@ IV-A Token-Level Features
Average Probability ($avg\_prob$): According to C2, real secret tokens generally have higher probabilities than fake ones. To capture this while addressing the instability of using a single token’s probability, we use the average probability of tokens selected in previous decoding steps as a feature. Compared with token’s probability sequence, the average probability could reduce computational complexity while ensuring interpretability (see our replication package [46]). In Fig. 5, the probabilities of tokens up to “2” are 0.99, 0.18, 0.03, 0.02, and 0.03, with an average of 0.25, which is set as the Average Probability for token “2”.
@@ IV-B Token Scoring Model Construction
Training Data Construction. We follow a process similar to our characterization study setup (Section III-B). We create completion prompts for each secret type based on searched code files and feed them into a Proxy Code LLM to generate candidate secrets. The generated candidates are validated and split into a real set $\mathcal{R}$ and a fake set $\mathcal{F}$. During the proxy Code LLM’s token decoding process, we extract the feature vector $\mathbf{feat}$ for each token, resulting in two vector sets: $\mathbf{Feat}_{\mathcal{R}}$ and $\mathbf{Feat}_{\mathcal{F}}$, corresponding to $\mathcal{R}$ and $\mathcal{F}$, respectively.
Scoring Model Training. To effectively distinguish between tokens of real and fake secrets, we train a Linear Discriminant Analysis (LDA) [48] model to find a combination of the four features that maximizes the differentiation between $\mathbf{Feat}_{\mathcal{R}}$ and $\mathbf{Feat}_{\mathcal{F}}$. We choose this linear model over complex models like Multi-Layer Perceptron (MLP) [49] because our feature vector for each token contains only four features, and using non-linear models can lead to overfitting issues with low-dimensional feature vectors.
We prioritize the resource-efficient LDA rather than other complex models due to the immediacy required in decoding.
@@ 11             $Beam\leftarrow Cands$
  • Masking Invalid Tokens (line 8). In traditional beam search, the LLM predicts the probability distribution $\mathbf{p}$ based on the existing token sequence $seq$ (line 7), and the next step is to select the top $B$ tokens from this distribution (line 9). However, in our secret extraction scenario, some tokens with unsupported characters are invalid (e.g., “*”), as secrets must adhere to specific formats. To address this and avoid invalid candidates for the scoring model, we mask invalid tokens in $\mathbf{p}$ by setting their probabilities to 0 (i.e. the MaskInvalidToks procedure), ensuring they are not selected in the subsequent step. The specific constraint for determining invalid tokens for each secret type is listed as a regular expression in Table I.
@@
    • Calling Scoring Model (line 12). Based on the extracted $\mathbf{feat}$, we call the trained scoring model to predict $prob_{\mathcal{R}}$ for $tok$ using Equation 1, indicating the probability that the given $tok$ belongs to a real secret under the condition that it is selected by LLM.
@@
Additional Optimization. Based on C1 from our characterization study, we optimize the beam search process (lines 18-20) by retaining all expanded candidate hypotheses ($Cands$) during the first $K$ steps instead of selecting the top $B$ hypotheses. This prevents missing real secrets due to the initially low probabilities of early tokens [50]. We set $K$ to 4 based on our observations in small-scale experiments and computational cost considerations.
@@ V Experimental Setup
  • RQ3.a: What are the effects of components of DeSec on secret extraction?
  • RQ3.b: How effective is our token scoring model in identifying secret tokens?
  • RQ4: How generalizable is DeSec with respect to different Code LLMs and types of secrets?
V-A Evaluation Dataset
@@
V-B Victim Code LLMs
@@ V-D Metrics
V-E Implementation
In the online decoding, the beam search size $B$ is set to 5, balancing token search breadth and text quality [54]. The hyper-parameter w for the LLM-predicted token probability is set to 0.75 by evaluating the results within interval [0,1]. Fig. 6 shows that it yields the highest number of real secrets and the best overall performance for generating both plausible secrets (PS#) and real secrets (RS#).
Note that Token Scoring Model Construction and Scoring Model Guided Decoding involve no randomness, ensuring reproducibility; thus, we do not conduct repeated experiments.
@@ VI Results and Analysis
Comparison to HCR. DeSec significantly outperforms HCR in extracting plausible secrets, with PRimprovements ranging from 16.8% to 51.2% across the five victim Code LLMs. DeSec outperforms HCR for most secret types, except for GOCI in DeepSeek-Coder-6.7B and CodeLlama-13B, where HCR extracts 1 and 9 more plausible secrets, respectively. The superiority of DeSec can be attributed to: (i) The excessively long context of HCR distracts the model’s attention[43], introducing noise and reducing focus on the target secret , and (ii) beam search decoding in DeSec being more effective than greedy search in HCR, evidenced by comparing HCR and BS-5 results.
@@
DeSec consistently demonstrates high effectiveness in extracting real secrets across all five victim Code LLMs, extracting between 95 and 358 real secrets, and reaches a total of 1076 (95+199+214+358+210), surpassing the numbers obtained by HCR (230 real secrets) and BS-5(845 real secrets).
The higher RS# for CodeLlama-13B may indicate that it memorizes more real secrets, posing more severe privacy leakage risks.
@@
VI-C RQ3.a: Ablation Study
To investigate the effects of the two key enhancements in DeSec, namely masking invalid tokens and scoring model-guided probability re-weighting step, we conduct an ablation study. We respectively remove the two enhancements, resulting in two variants of DeSec: w/o masking and w/o scoring.
@@
Answer to RQ3.a Token scoring model effectively prompts the models to output memorized secrets. Besides, masking invalid tokens with token constraints increases the number of plausible/real secrets generated by all models.
VI-D RQ3.b: Effectiveness in Identifying Secret Tokens
@@
Answer to RQ3.b The token scoring model demonstrates high values across various metrics for most models, indicating its effectiveness in identifying secret tokens.
VI-E RQ4: Generalizability Of DeSec
Generalizability across different models. Fig. 7(c) shows that, despite the scoring model being trained using StarCoder-15B as the proxy Code LLM, DeSec exhibits superior RS# for DeepSeek-Coder-6.7B (214) and CodeLlama-13B (358) compared to StarCoder-15B (210), suggesting that the identified characteristics (C1-C4) and the trained scoring model are generalizable to various Code LLMs.
However, there are great performance variations across models. Based on observations, we attribute this to intrinsic model characteristics which lead to the memory gap, like parameter scale, training strategy, and the frequency of secret text in the training data [55]. Specifically, DeSec relies on the model’s ability to memorize secrets during training. If a model has not memorized many secrets, DeSec cannot generate such secrets during decoding, as it cannot create new information that the model has not learned. Allowing the model to explore a broader range of possibilities during the decoding process, guided by DeSec’s scoring model, could be a potential mitigation approach to address this issue without altering the model’s memory capacity.
Generalizability to other secret types. To further validate the generalizability of DeSec to a wider range of secrets, we collect eight additional types of secrets, with 100 prompts for each type. We do not gather additional features for these secret types; instead, we directly utilize the token features collected in Section V-E to train the Token Scoring Model. Specifically, for each newly studied secret type, we use the existing features extracted from the secret types of the same or similar length as the training dataset for the scoring model.
Table VI and VII present the performance of two baselines and DeSec on these eight secret types in terms of the PS# and RS# metrics, respectively. DeSec achieves significant improvements in PS# across all types of secrets for all LLMs, with CodeLlama-13B reaching the highest value of 253. DeSec achieves RS# improvements on StableCode-3B, DeepSeekC-6.7B, and CodeLlama-13B. However, on StarCoder2-15B, due to the influence of paired key contexts (Stripe Live Secret Key and Stripe Test Secret Key), both BS-5 and DeSec performed worse than HCR. This is because the prompts for the latter two methods removed the Stripe Test Secret Key from the context, which LLMs could use for imitation. For the AWS Access Key ID, which appears most frequently on GitHub among the eight secret types, DeSec and the baselines only extract few real secrets. This is likely because the LLMs have predominantly memorized widely shared examples, such as “AKIAIOSFODNN7EXAMPLE”.
TABLE VI: PS# for extracting newly studied secrets
StableCode-3B CodeGen-7B DeepSeekC-6.7B CodeLlama-13B StarCoder2-15B
Secret Type HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec
Aws Access Key ID 0 2 4 0 7 7 0 5 9 0 15 17 1 1 2
Google Oauth Client Secret 0 0 38 0 4 49 3 16 48 0 21 66 0 2 73
Midtrans Sandbox Server Key 0 7 61 0 9 41 0 4 7 0 23 24 0 0 8
Flutterwave Live Api Secret Key 8 4 19 0 4 11 0 6 50 0 26 45 0 21 48
Flutterwave Test Api Secret Key 12 22 40 0 13 26 0 23 53 0 31 57 0 21 35
Stripe Live Secret Key 12 15 40 0 18 48 6 27 40 6 11 43 6 8 10
Ebay Production Client ID 0 0 0 0 0 0 0 0 0 1 0 0 2 3 19
Github Personal Access Token 0 3 34 0 3 50 1 3 2 0 0 1 0 0 14
Total 32 53 236 0 58 232 10 84 209 7 127 253 9 56 209
TABLE VII: RS# for extracting newly studied secrets
StableCode-3B CodeGen-7B DeepSeekC-6.7B CodeLlama-13B StarCoder2-15B
Secret Type HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec HCR BS-5 DeSec
Aws Access Key ID 0 1 1 0 4 4 0 3 3 0 3 3 0 0 0
Google Oauth Client Secret 0 0 0 0 0 0 0 0 0 0 24 23 0 0 0
Midtrans Sandbox Server Key 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Flutterwave Live Api Secret Key 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Flutterwave Test Api Secret Key 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Stripe Live Secret Key 1 1 7 0 1 1 2 1 14 1 2 24 2 1 1
Ebay Production Client ID 0 0 0 0 0 0 0 0 0 0 0 0 0 0 0
Github Personal Access Token 0 0 0 0 0 0 1 0 0 0 0 0 0 0 0
Total 1 2 8 0 5 5 3 4 17 1 29 50 2 1 1
Answer to RQ4 DeSec demonstrates strong generalizability across different models and effectively enhances the secret extraction effectiveness for other types of secrets.
VI-F Case Study
Fig. 8 illustrates two secret generation examples for DeSec, HCR, and BS-5. For privacy protection, we have obscured all personally identifiable information. We also provide other examples in our replication package [46].